# Graph Neural Networks for Job Shop Scheduling Problems: A Survey

### v1

@@ 2 Problem Description
Basic JSSP. A standard JSSP instance consists of a set of jobs $\mathcal{J}$ and a set of machines $\mathcal{M}$. Each job $J_{i}\in\mathcal{J}$ has an operation set containing $n_{i}$ operations $O_{i}=\{O_{ij}\}_{j=1}^{n_{i}}$ that must be processed in a specific order (i.e. precedence constraints), represented as $O_{i1}\to\dots\to O_{i{n_{i}}}$. Here, each $O_{ij}$ signifies an operation of $J_{i}$ with a processing time $p_{ij}\in\mathbb{N}$. Each machine can only process one job at a time, and preemption is not allowed. The objective in solving a JSSP instance is to determine a start time $S_{ij}$ for each operation $O_{ij}$ such that the makespan $C_{\max}=\max_{i,j}\{C_{ij}=S_{ij}+p_{ij}\}$ is minimized while adhering to all constraints. The size of a JSSP instance is $|\mathcal{J}|\times|\mathcal{M}|$.
@@
Figure 1: Disjunctive graph representation of basic JSSPs. Left panel represents a 3 (jobs) $\times$ 3 (machines) JSSP instance. The black arrows are conjunctive arcs, representing the precedence among operations within the same job. The dotted lines are disjunctive arcs whose directions are to be assigned. The disjunctive arcs (or the operation nodes) with the same color require the same machine for processing. Right panel represents a feasible solution. Best viewed in color.
Flexible JSSP. In JSSPs, each operation within a job must be processed on a specific machine. In contrast, the flexible job shop scheduling problem (FJSSP) allows each operation to be processed on any machine from a subset of available machines capable of performing that operation. Formally, each operation $O_{ij}$ can be processed on any machine $M_{k}$ from its compatible set $M_{k}\in\mathcal{M}_{ij}\subseteq\mathcal{M}$ with a processing time $p_{ijk}$. Each operation could be connected to multiple disjunctive arcs in the disjunctive graph representation, and thus, solving FJSSP is equivalent to selecting a disjunctive arc for each operation node and fixing its direction. An example of the disjunctive graph and a feasible solution of an FJSSP instance is illustrated in Fig. 2.
@@
Dynamic JSSP. Dynamic job shop scheduling problem (DyJSSP) adds a layer of complexity to the traditional JSSP by introducing elements of change and uncertainty that occur over time. Concretely, during the execution of the schedule, new jobs may arrive unpredictably, or uncertain events may happen, such as machine breakdowns, job cancellations, or variations in job processing times. This requires the schedule to be flexible and often necessitates real-time adjustments.
Distributed JSSP. Distributed job shop scheduling problem (DiJSSP) is another variant of the basic JSSP that incorporates spatial distribution among the machines involved. Formally, a DiJSSP instance comprises a set of factories =$\mathcal{F}=\{f_{i}\}_{i=1}^{|\mathcal{F}|}$, each equipped with its own machines. The problem involves a collection of jobs that must be processed in one factory. In DiJSSP, two key decisions should be made: firstly, selecting and assigning each job to a suitable factory, and secondly, scheduling the operations of the jobs within their assigned factory. Once a job begins processing in a selected factory, it cannot be transferred to another. The primary objective is to minimize the maximum makespan across all factories. An example of the disjunctive graph representation and a feasible solution of a DiJSSP instance is illustrated in Fig. 3.
FSPs. In the flow-shop scheduling problem (FSP), all jobs must be processed on the same set of machines in the same order, while the machine sequence is fixed for all jobs in JSSPs. A hybrid flow-shop scheduling problem (HFSP) is a generalization of FSP where each stage can have one or more parallel machines. Jobs must still go through each stage in a fixed sequence, but they can be processed on any machine within each stage. For permutation flow-shop scheduling problem (PFSP), not only must all jobs go through the machines in the same order, but the order of jobs must also be the same on all machines.
@@ 4.1 Non-DRL Methods
Corsini et al. (2024) propose a self-supervised learning method, called self-labeling pointer network (SPN) as an alternative to DRL, to create a solution construction agent for JSSPs. In this self-supervised training strategy, multiple parallel solutions are sampled for each problem instance. Then, the best solution is used to provide pseudo-labels, and the cross-entropy loss using these pseudo-labels is minimized. A pointer network using a GAT encoder and a multi-head attention decoder is trained through this method. In specific, the model is trained using instances generated through Taillard’s method (Taillard, 1993) with sizes of $10\times 10$, $15\times 10$, $15\times 15$, $20\times 10$, $20\times 15$, and $20\times 20$. The method is evaluated on the TA and DMU (Demirkol et al., 1998) benchmark datasets. Their method outperforms priority dispatching rules (PDRs), the insertion algorithm (Nowicki and Smutnicki, 1996), learning to dispatch (L2D)  (Zhang et al., 2020), the curriculum learning method (Iklassov et al., 2022), and the enhanced local search metaheuristic (Falkner et al., 2022) by a clear margin. Especially when using the sampling inference method, optimality gaps compared to L2D are more than half lower. The performance seems to match, or in some instances outperform, the best DRL-based methods for the JSSP. Thus, this method may offer a promising direction for further research. However, due to the autoregressive inference method without a Markov decision process (MDP) formulation, it may be somewhat limited to static problems. For dynamic problems, a purely autoregressive method may not be suitable without more advanced adaptations. More recently, Echeverria et al. (2024) developed a heterogeneous GAT to compute the embeddings of edges between machines and operations, and then an MLP with Softmax function is used to estimate the probabilities of edge selection, meaning the assignment of operations to machines. The heterogeneous GAT is trained by behavioral cloning (BC), with the optimal solutions as supervised signals. Table 1 provides a summary of GNN-based methods for solving JSSP without using (D)RL.
@@
Echeverria et al. (2024) 2024 FJSSP Solution Construction Heterogeneous Graph Heterogeneous GAT
@@ 4.2.1 Preliminaries on DRL
DRL combines traditional RL with deep learning techniques to effectively tackle sequential decision-making problems. Given an MDP formulation of a sequential decision-making problem, an agent (parameterized by a policy network $\pi_{\theta}$) sequentially takes actions according to the varying states derived from the environment. The environment iteratively responds to the actions by providing rewards to the agent and transferring the states. The objective of DRL is to optimize the policy network such that the expected reward is maximized, which is usually achieved by learning an optimal policy $\pi^{*}_{\theta}$, or an optimal state-action value network $Q^{*}_{\delta}$ so that the optimal policy can be retrieved later from $Q^{*}_{\delta}$, that is, $\pi^{*}(a|s)=\arg\max_{a}Q^{*}_{\delta}(s,a)$. Generally, DRL algorithms are classified into policy-based methods, including REINFORCE, A3C, and PPO (Schulman et al., 2017; Williams, 1992; Mnih et al., 2016), and value-based methods, including DQN, double DQN, and dueling DQN (Wang et al., 2016; Mnih et al., 2015; Van Hasselt et al., 2016). We refer interested readers to François-Lavet et al. (2018); Tassel et al. (2021) for more knowledge on DRL.
@@ 4.2.3 Basic JSSPs
While most research primarily focuses on constructive methods, Falkner et al. (2022) and Zhang et al. (2024) aim to improve search methods. Falkner et al. (2022) propose NeuroLS, a DRL approach designed to steer local search for combinatorial optimization. They integrate a GNN to capture information on current solutions found by the local search engine, combining it with state information from the local search itself. They train their model using a DDQN approach on instances generated by Taillard’s method, ranging in sizes from $15\times 15$ to $30\times 20$. The evaluation is conducted on the TA benchmark dataset. With 100 iterations of the local search, NeuroLS outperforms L2D, GNNRL, and ScheduleNet, as well as traditional optimization methods like simulated annealing, iterated local search, and variable neighborhood search (VNS) on most problem sizes, though the improvement over VNS is limited. Zhang et al. (2024) opt to use a GNN-based DRL approach to create an improvement heuristic. In their method, the agent guides a neighborhood-based improvement heuristic by selecting node-pairs to swap within the disjunctive graph. They train their method on synthetic instances generated using Taillard’s method, with sizes of $10\times 10$, $15\times 10$, $15\times 15$, $20\times 10$, and $20\times 15$. The evaluation encompasses the TA, ABZ, FT, LA, SWV, ORB, and YN benchmarks, along with three very large datasets containing up to 1000 jobs. Although this method increases runtime, it significantly outperforms construction-based DRL methods on the benchmarks, achieving near-optimal solutions for some instances. Furthermore, the method proves competitive against a tabu search approach while being faster. Notably, on the large datasets, within a few minutes of computation, the method achieves 15 to 25 percent improvements over solutions obtained using CP-SAT with a 1-hour time limit. This method outlines a promising new research direction for the application of DRL to improvement heuristics in scheduling.
@@
Echeverria et al. (2023) 2023 FJSSP Dispatching Rule Generation PPO Heterogeneous Graph Heterogeneous GAT
@@ 4.2.4 Flexible JSSPs
In consideration of system complexities, GNNs have increasingly been employed to synergize feature extraction with reinforcement learning techniques. Zeng et al. (2022b) formulate the FJSSP as an MDP where the disjunctive graph is used to represent the state, and operation set and machine allocation are used as the actions. A powerful GIN is used to extract the state features. The asynchronous advantage actor-critic (A3C) algorithm is selected for optimization training, with performance evidenced on Brandimarte instances. Dax et al. (2022) use Q-Learning to train GNNs for tackling the FJSSP. They utilize disjunctive graphs to simulate states and encode an action value function with heterogeneous GCNs. Benchmark testing shows the algorithm’s advantages in terms of training time and solution quality compared to metaheuristic algorithms. Song et al. (2022) propose an end-to-end DRL algorithm to learn PDRs for the FJSSP. An MDP formulation integrates the operation selection and machine assignment into a single decision. They expand the disjunctive graph with machine nodes and introduce a heterogeneous graph to describe the state. The PPO algorithm is applied to train the policy network. Experiments show their approach outperforms traditional PDRs. Zhang and Li (2023) propose a hierarchical graph structure model that synergistically integrates the environment and decision-making module. They formulate the environment module as a heterogeneous graph to define multi-relational FJSSP instances and use a hierarchical GNN. The decision module incorporates actor-critic networks to train reasoning strategies and enhance decision performance. This approach outperforms PDRs and DRL-based methods in Brandimarte and Hurink test cases. Echeverria et al. (2023) formulate the FJSSP as an MDP by adding a new job-type node and removing irrelevant nodes and edges, significantly shrinking the action space and enhancing the policy’s ability to capture information. They train their model using PPO. Their approach outperforms dispatching rules on a variety of Brandimarte instances, particularly large ones. Farahani et al. (2023) propose an edge feature guided relational graph attention-based deep reinforcement learning (ERGAT-DRL) framework to address the FJSSP considering the sequence of setup times. A weighted relational graph is employed to address the machine flexibility and sequence dependency. They utilize an advantage actor-critic (A2C) algorithm to train their model. Their method outperforms PDRs on realistic data. Wang et al. (2023a) formulate the FJSSP using a tight state representation for describing operations and machines. Additionally, an end-to-end learning framework, which combines self-attention models with DRL, is introduced. A dual attention network (DAN) for extracting features between operations and machines is trained using PPO in the training phase. Experiments on synthetic and public benchmark datasets from Brandimarte and Hurink have confirmed the superiority over traditional PDRs and state-of-the-art DRL methods. Lei et al. (2022b) introduce a multi-pointer graph network (MPGN) architecture to embed local states effectively. This architecture delineates two sub-policies: job operation action and machine action. During the training phase, a multi-PPO approach is adopted to learn these sub-policies. Benchmark results show that the proposed method achieves superior generalization performance.
@@ 5 Discussion
This section summarizes the main findings of our review. The existing GNN-based methods rely on standard GNN types such as GCN, GAT, and GIN. Additionally, heterogeneous GNNs are widely applied to better represent jobs, machines, and operations, to enhance performance. On the other hand, most GNN-based methods are developed with RL algorithms since achieving high-quality solutions as labels in supervised learning is computationally expensive, especially when the problems become more complicated and large-scale. The REINFORCE and PPO algorithms are mostly used in literature, as they have shown promising results in several early attempts. We provide a detailed summary for each scheduling problem below.

### v2

@@ 2 Problem Description
Open-shop Scheduling Problem. In open-shop scheduling problem (OSP), all jobs must be processed on a set of machines, with the objective of minimizing the total time to complete all jobs, known as makespan. There are no predefined orders in which jobs must visit the machines, and each job must be processed once on each machine.
Flow-shop Scheduling Problem. Building upon the OSP, the flow-shop scheduling problem (FSP) requires that each job is processed on a set of machines, while the machine sequence is the same and fixed for all jobs. A hybrid flow-shop scheduling problem (HFSP) is a generalization of FSP where each stage can have one or more parallel machines. Jobs must still go through each stage in a fixed sequence, but they can be processed on any machine within each stage. For permutation flow-shop scheduling problem (PFSP), not only must all jobs go through the machines in the same order, but the order of jobs must also be the same on all machines.
Figure 1: Disjunctive graph representation of JSSPs. Left panel represents a 3 (jobs) $\times$ 3 (machines) JSSP instance. The black arrows are conjunctive arcs, representing the precedence among operations within the same job. The dotted lines are disjunctive arcs whose directions are to be assigned. The disjunctive arcs (or the operation nodes) with the same color require the same machine for processing. Right panel represents a feasible solution. Best viewed in color.
Job-shop Scheduling Problem. Unlike FSPs, each job has a unique sequence of operations that must be processed on specific machines in the job-shop scheduling problem (JSSP). Formally, a standard JSSP instance consists of a set of jobs $\mathcal{J}$ and a set of machines $\mathcal{M}$. Each job $J_{i}\in\mathcal{J}$ has an operation set containing $n_{i}$ operations $O_{i}=\{O_{ij}\}_{j=1}^{n_{i}}$ that must be processed in a specific order (i.e. precedence constraints), represented as $O_{i1}\to\dots\to O_{i{n_{i}}}$. Here, each $O_{ij}$ signifies an operation of $J_{i}$ with a processing time $p_{ij}\in\mathbb{N}$. Each machine can only process one job at a time, and preemption is not allowed. The objective in solving a JSSP instance is to determine a start time $S_{ij}$ for each operation $O_{ij}$ such that the makespan $C_{\max}=\max_{i,j}\{C_{ij}=S_{ij}+p_{ij}\}$ is minimized while adhering to all constraints. The size of a JSSP instance is $|\mathcal{J}|\times|\mathcal{M}|$.
@@
Distributed Job-shop Scheduling Problem. Distributed job shop scheduling problem (DiJSSP) is a variant of the basic JSSP that incorporates spatial distribution among the machines involved. Formally, a DiJSSP instance comprises a set of factories =$\mathcal{F}=\{f_{i}\}_{i=1}^{|\mathcal{F}|}$, each equipped with its own machines. The problem involves a collection of jobs that must be processed in one factory. In DiJSSP, two key decisions should be made: firstly, selecting and assigning each job to a suitable factory, and secondly, scheduling the operations of the jobs within their assigned factory. Once a job begins processing in a selected factory, it cannot be transferred to another. The primary objective is to minimize the maximum makespan across all factories. An example of the disjunctive graph representation and a feasible solution of a DiJSSP instance is illustrated in Fig. 2.
Flexible Job-shop Scheduling Problem. Different from JSSPs, the flexible job shop scheduling problem (FJSSP) allows each operation to be processed on any machine from a subset of available machines capable of performing that operation. Formally, each operation $O_{ij}$ can be processed on any machine $M_{k}$ from its compatible set $M_{k}\in\mathcal{M}_{ij}\subseteq\mathcal{M}$ with a processing time $p_{ijk}$. Each operation could be connected to multiple disjunctive arcs in the disjunctive graph representation, and thus, solving FJSSP is equivalent to selecting a disjunctive arc for each operation node and fixing its direction. An example of the disjunctive graph and a feasible solution of an FJSSP instance is illustrated in Fig. 3.
Dynamic Job-shop Scheduling Problem. Dynamic job shop scheduling problem (DyJSSP) adds a layer of complexity to the traditional JSSP by introducing elements of change and uncertainty that occur over time. Concretely, during the execution of the schedule, new jobs may arrive unpredictably, or uncertain events may happen, such as machine breakdowns, job cancellations, or variations in job processing times. This requires the schedule to be flexible and often necessitates real-time adjustments.
Dynamic Flexible Job-shop Scheduling Problem. By incorporating both flexibility in job routing from FJSSP and dynamic changes in the production environment from DyJSSP, the dynamic flexible job-shop scheduling problem (DyFJSSP) offers a realistic model for optimizing production in complex, dynamic environments. Its ability to respond to real-time changes makes it suitable for modern industries that require high levels of efficiency and adaptability, such as automotive assembly lines, semiconductor manufacturing, and custom product fabrication.
In addition to directed acyclic graphs, there is a recent interest in leveraging alternative graph-based techniques, such as Petri nets, to represent scheduling problems in a more dynamic and concurrent framework. One advantage of Petri nets is their direct interaction with deep reinforcement learning (DRL) agents. Unlike disjunctive graphs, which require additional methods to extract features for optimization, Petri nets naturally integrate with DRL by allowing token distributions to serve as part of the system’s observation space. This capability makes Petri nets a promising approach for solving scheduling problems (Lassoued and Schwung, 2024).
@@ 4.1 Non-DRL Methods
Corsini et al. (2024) propose a self-supervised learning method, called self-labeling pointer network (SPN) as an alternative to DRL, to create a solution construction agent for JSSPs. In this self-supervised training strategy, multiple parallel solutions are sampled for each problem instance. Then, the best solution is used to provide pseudo-labels, and the cross-entropy loss using these pseudo-labels is minimized. A pointer network using a GAT encoder and a multi-head attention decoder is trained through this method. In specific, the model is trained using instances generated through Taillard’s method (Taillard, 1993) with sizes of $10\times 10$, $15\times 10$, $15\times 15$, $20\times 10$, $20\times 15$, and $20\times 20$. The method is evaluated on the TA and DMU (Demirkol et al., 1998) benchmark datasets. Their method outperforms priority dispatching rules (PDRs), the insertion algorithm (Nowicki and Smutnicki, 1996), learning to dispatch (L2D)  (Zhang et al., 2020), the curriculum learning method (Iklassov et al., 2022), and the enhanced local search metaheuristic (Falkner et al., 2022) by a clear margin. Especially when using the sampling inference method, optimality gaps compared to L2D are more than half lower. The performance seems to match, or in some instances outperform, the best DRL-based methods for the JSSP. Thus, this method may offer a promising direction for further research. However, due to the autoregressive inference method without a Markov decision process (MDP) formulation, it may be somewhat limited to static problems. For dynamic problems, a purely autoregressive method may not be suitable without more advanced adaptations. Table 1 provides a summary of GNN-based methods for solving JSSP without using (D)RL.
@@ 4.2.1 Preliminaries on DRL
DRL combines traditional RL with deep learning techniques to effectively tackle sequential decision-making problems. Given an MDP formulation of a sequential decision-making problem, an agent sequentially takes actions according to the varying states derived from the environment. The environment iteratively responds to the agent’s actions by providing rewards and transitioning to new states. The objective of DRL is to find an optimal policy $\pi^{*}_{\theta}$ that maximized the expected cumulative reward (i.e., return). Generally, policy-based and value-based approaches are two primary methods for learning optimal policies. Policy-based approaches, such as REINFORCE, A3C, and PPO (Schulman et al., 2017; Williams, 1992; Mnih et al., 2016), directly learns the policy $\pi_{\theta}$, which defines the probability distribution over actions for each state. In contrast, value-based approaches, such as DQN, double DQN, and dueling DQN (Wang et al., 2016; Mnih et al., 2015; Van Hasselt et al., 2016), focuses on learning value functions $Q^{*}_{\delta}$, which estimate the expected cumulative reward for each state (or state-action pair). We refer interested readers to François-Lavet et al. (2018); Tassel et al. (2021) for more knowledge on DRL.
@@ 4.2.2 MDP formulations of JSSPs
van Remmerden et al. (2024) 2024 JSSP Dispatching Rule Generation Offline RL with Conservative Qlearning (CQL) Disjunctive Graph GIN
@@ 4.2.3 Basic JSSPs
While most research primarily focuses on constructive methods, Falkner et al. (2022) and Zhang et al. (2024) aim to improve search methods. Falkner et al. (2022) propose NeuroLS, a DRL approach designed to steer local search for combinatorial optimization. They integrate a GNN to capture information on current solutions found by the local search engine, combining it with state information from the local search itself. They train their model using a DDQN approach on instances generated by Taillard’s method, ranging in sizes from $15\times 15$ to $30\times 20$. The evaluation is conducted on the TA benchmark dataset. With 100 iterations of the local search, NeuroLS outperforms L2D, GNNRL, and ScheduleNet, as well as traditional optimization methods like simulated annealing, iterated local search, and variable neighborhood search (VNS) on most problem sizes, though the improvement over VNS is limited. Zhang et al. (2024) opt to use a GNN-based DRL approach to create an improvement heuristic. In their method, the agent guides a neighborhood-based improvement heuristic by selecting node-pairs to swap within the disjunctive graph. They train their method on synthetic instances generated using Taillard’s method, with sizes of $10\times 10$, $15\times 10$, $15\times 15$, $20\times 10$, and $20\times 15$. The evaluation encompasses the TA, ABZ, FT, LA, SWV, ORB, and YN benchmarks, along with three very large datasets containing up to 1000 jobs. Although this method increases runtime, it significantly outperforms construction-based DRL methods on the benchmarks, achieving near-optimal solutions for some instances. Furthermore, the method proves competitive against a tabu search approach while being faster. Notably, on large datasets, within a few minutes of computation, the method achieves 15 to 25 percent improvements over solutions obtained using CP-SAT with a 1-hour time limit. This method outlines a promising new research direction for the application of DRL to improvement heuristics in scheduling.
Most existing DRL approaches for JSSPs are online RL, where an RL agent selects actions by interacting with an environment. One downside of online RL is that it cannot learn from existing data, for instance, solutions generated from optimization algorithms, although such data is easily obtained as many fast heuristics exist. Imitation learning utilizes existing data but is hard to obtain decisions better than the algorithm used to generate training data. van Remmerden et al. (2024) propose the first Offline RL method to solve JSSP, Offline Reinforcement Learning for Learning to Dispatch (Offline-LD), which makes use of existing data and overcomes the limitation of imitation learning. Offline-LD is developed based on two CQL-based Q-learning methods (mQRDQN and discrete mSAC), and utilizes Conservative Qlearning (CQL) to enable d-mSAC and mQRDQN to learn from existing datasets. They show that with Offline-LD, the existing online RL approach, such as L2D Zhang et al. (2020), can be easily adapted as an offline RL method. The experiments demonstrate that Offline-LD, trained with only 100 instances, outperforms online RL (L2D), behavioral cloning approach, and dispatching rules on generated and benchmark instances. In addition, the results show that less perfect data, generated with noises, can achieve similar or better policies than those obtained from the expert dataset generated by a CP solver. This work provides a new promising RL approach to solving JSSP and its variants.
@@ 4.2.4 Flexible JSSPs
In consideration of system complexities, GNNs have increasingly been employed to synergize feature extraction with reinforcement learning techniques. Zeng et al. (2022b) formulate the FJSSP as an MDP where the disjunctive graph is used to represent the state, and operation set and machine allocation are used as the actions. A powerful GIN is used to extract the state features. The asynchronous advantage actor-critic (A3C) algorithm is selected for optimization training, with performance evidenced on Brandimarte instances. Dax et al. (2022) use Q-Learning to train GNNs for tackling the FJSSP. They utilize disjunctive graphs to simulate states and encode an action value function with heterogeneous GCNs. Benchmark testing shows the algorithm’s advantages in terms of training time and solution quality compared to metaheuristic algorithms. Song et al. (2022) propose an end-to-end DRL algorithm to learn PDRs for the FJSSP. An MDP formulation integrates the operation selection and machine assignment into a single decision. They expand the disjunctive graph with machine nodes and introduce a heterogeneous graph to describe the state. The PPO algorithm is applied to train the policy network. Experiments show their approach outperforms traditional PDRs. Zhang and Li (2023) propose a hierarchical graph structure model that synergistically integrates the environment and decision-making module. They formulate the environment module as a heterogeneous graph to define multi-relational FJSSP instances and use a hierarchical GNN. The decision module incorporates actor-critic networks to train reasoning strategies and enhance decision performance. This approach outperforms PDRs and DRL-based methods in Brandimarte and Hurink test cases. Farahani et al. (2023) propose an edge feature guided relational graph attention-based deep reinforcement learning (ERGAT-DRL) framework to address the FJSSP considering the sequence of setup times. A weighted relational graph is employed to address the machine flexibility and sequence dependency. They utilize an advantage actor-critic (A2C) algorithm to train their model. Their method outperforms PDRs on realistic data. Wang et al. (2023a) formulate the FJSSP using a tight state representation for describing operations and machines. Additionally, an end-to-end learning framework, which combines self-attention models with DRL, is introduced. A dual attention network (DAN) for extracting features between operations and machines is trained using PPO in the training phase. Experiments on synthetic and public benchmark datasets from Brandimarte and Hurink have confirmed the superiority over traditional PDRs and state-of-the-art DRL methods. Lei et al. (2022b) introduce a multi-pointer graph network (MPGN) architecture to embed local states effectively. This architecture delineates two sub-policies: job operation action and machine action. During the training phase, a multi-PPO approach is adopted to learn these sub-policies. Benchmark results show that the proposed method achieves superior generalization performance.
@@ 5 Discussion
This section summarizes the main findings of our review. Before delving into the specific characteristics of each scheduling variant, we first address several recurring challenges that reoccur across all subtypes of scheduling problems. These include issues such as the lack of consensus on representation choices, the absence of standardized benchmarking, and concerns related to evaluation metrics, scalability, and adaptability.
5.1 Recurring Challenges:
  • 1.
    Representation Consensus: For all problem variants—JSSPs, FJSSPs, DyJSSPs, DiJSSPs, and FSPs—there is no agreement on the optimal graph representation for learning. While disjunctive, heterogeneous, and bipartite graphs are commonly used, the lack of comparative studies prevents definitive conclusions about the most suitable format for various contexts. Furthermore, it’s plausible that there may be a better methodology for approaching these problems than graph representation. Alternative data structures or modelling techniques offer more efficient or insightful solutions. However, this area remains largely unexplored and could benefit from further research to ascertain the most effective approaches.
  • 2.
    Absence of Standardized Benchmarking: The lack of standardized benchmarking across different works poses another challenge. Many works use different problem instances, including synthetic or use-case-specific datasets, making it difficult to directly compare the performance of various approaches. Commonly used benchmarks in these studies are often outdated and fail to reflect the scale and complexity of contemporary industrial problems. This discrepancy hinders accurate assessments of these methods’ real-world applicability.
  • 3.
    Lack of Standardized Evaluation Metrics: Evaluation metrics differ between studies, with some focusing on optimality gaps in terms of makespan, while others consider one or more other objectives. This inconsistency in evaluation metrics adds to the difficulty of comparing methods across different problem variants.
  • 4.
    Scalability Issues: While some methods achieve impressive results, there is limited research into how well these methods scale as problem size and complexity increase. Few studies address the trade-offs between the quality of the solutions and computational efficiency for large instances. More work is needed to evaluate the scalability of GNN-based methods and their ability to handle real-world, large-scale scheduling problems.
5.2 Problem-Specific Observations:
  • 1.